# Tianxiang Chen USTC, China Alibaba Cloud, China

### v3

@@ Abstract
Keywords Mamba-in-Mamba  $\cdot$ State Space Model  $\cdot$ Infrared Small Target Detection
1 Introduction
@@
In recent years, deep-learning-based methods have significantly improved ISTD performances, and most of them are CNN-based methods [12, 13, 14, 15, 16, 17, 10, 18]. However, the drawback of CNN-based methods is that their focus on local features is at the cost of global contexts. Global contexts are important to ISTD because the background pixels and small targets often seem so similar that cannot be distinguished by local features alone, which easily triggers missed detection. As a solution, some hybrid methods [11, 19, 20, 21, 22, 23, 24] that combine ViT with CNN have been proposed to rely on ViT’s ability to model long-range dependencies. However, these methods generally suffer from a heavier computational burden due to the quadratic computational complexity of ViT. Despite certain work [19] adopting linear ViTs, its accuracy is still subordinate to the designs with quadratic complexity. Considering that high-resolution images are not rare in the infrared remote sensing domain (e.g. images produced by high-resolution infrared military sensors), this efficiency defect will be amplified when the resolution gets larger and hinder real-time ISTD. How to relieve the inefficiency while maintaining high ISTD accuracy is our main concern.
@@
We aim to tailor a pure Mamba-based encoder to solve the locality defect while still maintaining a superior model efficiency. To this end, we get inspired by TNT [30], which effectively models local structural features with a trivial increase of computation and memory cost, and propose a novel Mamba-in-Mamba (MiM-ISTD) architecture for more efficient ISTD. To boost the feature representation ability of Mamba, we divide the input image into several patches as "visual sentences" and then further divide them into sub-patches as "visual words". We use an outer Mamba block to extract features of visual sentences, and further assist it with inner Mamba blocks to excavate the local features of smaller visual words. In particular, features and relations between visual words in each visual sentence are calculated independently using a shared network to ensure the added amount of parameters and FLOPs is minimal. Then, the visual word features are consolidated back into their respective sentences. In this way, MiM-ISTD enables us to extract visual information with refined granularity. We present in Figure  1 that MiM-ISTD exhibits superior efficiency over other methods in terms of GPU memory usage and inference time, particularly as the resolution of infrared images increases. In general, MiM-ISTD can achieve the most notable accuracy-efficiency balance compared with other SOTA methods.
@@
2 Related Work
2.1 Infrared Small Target Detection Networks
@@
Generally speaking, present ISTD networks can be classified to two categories: CNN-based and hybrid networks. CNN-based networks mainly focus on local feature extraction. Dai et al. [16] propose asymmetric contextual modulation (ACM) for cross-layer information exchange to improve ISTD performance. They also design AlcNet [14], including a local attention module and a cross-layer fusion module to preserve the local features of small targets. Wang et al. propose MDvsFA [17], which applies generative adversarial network (GAN) to ISTD, and achieves a trade-off between miss detection and false alarm. BAUENet [13] introduces uncertainty to ISTD and achieves boundary-aware segmentation. DNANet [10] progressively interacts with high and low-level features. Dim2Clear [31] treats ISTD as an image detail reconstruction task by exploring the image enhancement idea. FC3-Net [32] explores feature compensation and cross-level correlation for ISTD task. ISNet [18] designs a simple Taylor finite difference-inspired block and a two-orientation attention aggregation module to detect targets.
@@
Relying solely on local features for ISTD may lead to missed detection of small targets, which can merge into similar backgrounds and become indistinguishable. Therefore, hybrid methods complement local details with global contexts by combining ViT with CNN. IRSTFormer[33] adopts hierarchical ViT to model long-range dependencies but lays insufficient emphasis on mining local details. ABMNet [19] adopts ODE methods in both CNN and linear ViT structure design for ISTD. IAANet [20] concatenates local patch outputs from a simple CNN with the original transformer, but causes limited feature extraction, especially in low-contrast scenarios. RKformer [24] applies the Runge-Kutta method to build coupled CNN-Transformer blocks to highlight infrared small targets and suppress background interference. However, most of these hybrid methods suffer from a quadratic computational complexity due to the usage of ViT. Despite some work [19] adopting linear ViT design, its detection accuracy cannot be on par with other works with quadratic-complexity ViTs.
@@
To improve network efficiency while maintaining high accuracy, we draw inspiration from both Mamba [25] and TNT [30], and propose a Mamba-in-Mamba architecture for ISTD, within which contains a pure Mamba-in-Mamba pyramid encoder that implements efficient local and global feature extraction.
2.2 Mamba in Vision Tasks
@@
Recently, state space sequence models (SSMs) [26] have shown promise in efficiently handling long sequence modeling, offering an alternative for addressing long-range dependencies in visual tasks. Compared with transformers, SSMs are more efficient since they scale linearly with sequence length, and retain a superior ability to model long-range dependencies. Several latest studies have demonstrated the effectiveness of Mamba in vision tasks. For instance, Vim [34] proposed a generic vision backbones with bidirectional Mamba blocks. VMamba [28] proposed a hierarchical Mamba-based vision backbone and a cross-scan module to address the issue of direction-sensitivity arising from the disparities between 1D sequences and 2D image representations.
@@
Notably, Mamba has been most widely applied to the medical image segmentation areas. U-Mamba [27], Vm-unet [35], Mamba-unet [36] and SegMamba [37] proposed a task-specific architecture with the Mamba block based on nnUNet [38], Swin-UNet [39], VMamba [28] and Swin-UNETR [40], respectively. P-Mamba [29] combined PM diffusion with Mamba to efficiently remove background noise while preserving target edge details. Swin-UMamba [41] verified that ImageNet-based pre-training is important to medical image segmentation for Mamba-based networks. Weak-Mamba-UNet [42] and Semi-Mamba-UNet [43] are for weak-supervised and semi-supervised medical image segmentation. Vivim [44] introduced Mamba for medical video object segmentation.
@@
Figure 2: Illustration of the 2D Selective Scan (SS2D) on an infrared image. We commence by scanning an image using scan expanding. The four resulting feature sequences are then individually processed through the S6 block, and the four output sequences are merged (scan merging) to construct the final 2D feature map.
@@
Since these Mamba models have achieved promising results in various vision tasks, we intend to study whether Mamba can also bring advancements to ISTD as it has brought to other vision tasks. However, when we directly apply visual Mamba blocks to ISTD, the detection accuracy is not considerable, despite superior model efficiency. The reason is that the targets in ISTD tasks are very small, which requires to pay more attention to the local features than other vision tasks with mainly common-size targets, while the original visual Mamba block cannot well explore these local features. To solve this defect, we propose Mamba-in-Mamba (MiM-ISTD) for ISTD, which takes visual sentences and visual words flows simultaneously and send each flows to respective visual Mamba blocks to obtain both local and global features with high efficiency.
3 Method
@@
3.1 Preliminaries
3.1.1 State Space Models.
State Space Models (SSMs) are commonly employed as linear time-invariant systems that transform a one-dimensional input stimulus $x(t)\in\mathbb{R}^{L}$ through intermediary implicit states $h(t)\in\mathbb{R}^{N}$ to an output $y(t)\in\mathbb{R}^{L}$. In mathematical terms, SSMs are typically described by linear ordinary differential equations (ODEs) (Equation  1), where the system is characterized by a set of parameters including the state transition matrix $A\in\mathbb{C}^{N\times N}$, the projection parameters $B,C\in\mathbb{C}^{N}$, and the skip connection $D\in\mathbb{C}^{1}$.
@@ 3.1 Preliminaries
3.1.2 Discretization.
State Space Models (SSMs) present significant challenges when applied to deep learning scenarios due to their continuous-time nature. To address this, the ODE needs to be transformed into a discrete function. Considering the input $x_{k}\in\mathbb{R}^{L\times D}$, a sampled vector within the signal flow of length $L$ following [45], a timescale parameter $\Delta$ can be introduced to the continuous parameters $A$ and $B$ into their discrete counterparts $\overline{A}$ and $\overline{B}$ following the zeroth-order hold (ZOH) rule. Consequently, Equation  1 can be discretized as follows:
$\begin{split}&h_{k}=\overline{A}h_{k-1}+\overline{B}x_{k}\\ &y(t)=Ch(t)+Dx(t)\\ &\overline{A}=e^{\Delta A}\\ &\overline{B}=(e^{\Delta A}-I)A^{-1}B\\ &\overline{C}=C\\ \end{split}$ (2)
@@
3.1.3 2D Selective Scan.
2D spatial information cannot be effectively captured by models designed for 1D data, making it unsuitable to directly apply Mamba to vision tasks. The 2D selective scan (SS2D) in [28] can address the issue. The overview of SS2D is depicted in Figure  2. SS2D arranges image patches in four different directions to generate four separate sequences. The quad-directional scanning strategy ensures that each element in the feature map integrates information from all other locations in different directions, thus creating a global receptive field without increasing linear computational complexity. Each resulting feature sequence is then processed using the selective scan space state sequential model (S6) before merging to reconstruct the 2D feature map. Given the input feature $z$, then the output feature $\overline{z}$ of SS2D can be expressed as:
@@
where $i\in\{1,2,3,4\}$ represents one of the four scanning directions. $expand(·)$ and $merge(·)$ refer to the scan expanding and scan merging operations in [28]. The S6 block in Equation  4 is the core SSM operator of SS2D. It enables each element in a 1D array to engage with any previously scanned samples through a condensed hidden state. For a more comprehensive understanding of S6, [28] provides an in-depth explanation.
3.2 Mamba-in-Mamba for Efficient ISTD
Present ISTD methods mainly typically employ CNNs or hybrids of CNNs and ViTs. The latter compensates for the shortcomings of the former in modeling long-range dependencies but suffers from a quadratic computational complexity. Recently, Mamba has been proposed. It is renowned for its superior model efficiency, less GPU memory usage, and better long-range dependency modeling and has been successfully applied to the vision domain. Therefore, we explore whether Mamba can also be applied to improve ISTD performances. However, when we directly apply visual mamba block to ISTD, the accuracy is not very impressive because local features, which matter a lot to detect small targets, are less explored. To address this deficiency, we propose a Mamba-in-Mamba (MiM-ISTD) architecture, shown in Figure  3, to learn both global and local information in an image while guaranteeing superb model efficiency.
@@ 3.2 Mamba-in-Mamba for Efficient ISTD
3.2.1 Convolutional Stem.
We construct a convolutional stem, where a stack of $3\times 3$ convolutions is utilized, to produce visual words $\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$ and visual sentences $\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times D}$ at the first stage, where $C$ is the visual word dimension and $D$ is the visual sentence dimension. Each visual word corresponds to a $2\times 2$ pixel region in the original image, and each visual sentence is composed of $4\times 4$ visual words at this stage. Unlike ViTs, we do not add the position embedding bias to visual words and sentences due to the causal nature of visual mamba block [28].
3.2.2 MiM Hierarchical Encoder.
The core part of our MiM is its hierarchical encoder of four stages with different numbers of tokens, as shown in Figure  3. Across the four stages, the spatial shape of visual words is set as $\frac{H}{2}\times\frac{W}{2}$, $\frac{H}{4}\times\frac{W}{4}$, $\frac{H}{8}\times\frac{W}{8}$ and $\frac{H}{16}\times\frac{W}{16}$. The spatial shape of visual sentences are set as $\frac{H}{8}\times\frac{W}{8}$, $\frac{H}{16}\times\frac{W}{16}$, $\frac{H}{32}\times\frac{W}{32}$ and $\frac{H}{64}\times\frac{W}{64}$. We adopt the patch merging in [39] as the down-sampling operation. Each stage consists of multiple MiM blocks which process both word-level and sentence-level features. The visual words $x^{i,j}$ are mapped to a sequence of word embeddings $w^{i,j}$ via a linear projection, expressed as:
@@
Figure 4: Overview of (a) our proposed Mamba-in-Mamba (MiM) block, which contains an inner Mamba block and an outer Mamba block, and (b) the visual Mamba block [28]. The inner Mamba block is shared in the same layer. The dashed line in (a) means bypassing the outer Mamba block.
The MiM block handles two different data streams: one traverses through the visual sentences, while the other manages the visual words within each sentence. For the word embeddings, MiM block explores the relation between visual words:
$W^{i}_{l}=W^{i}_{l-1}+Mamba(LN(W^{i}_{l-1})),l=1,2,...,L$ (7)
where $l$ is the index of the $l-th$ block, and $L$ is the total count of such blocks. The first word embedding input $W^{i}_{0}$ is the $W^{i}$ in Equation  6. All transformed word embeddings are denoted by $\mathcal{W}_{l}=[W^{1}_{l},W^{2}_{l},...,W^{n}_{l}]$. This can be viewed as an inner Mamba block, illustrated in Figure  4 (a). In this process, the relationships among visual words are built by computing interactions among each two visual words. For example, in a patch containing a small target, a word denoting the target would have a stronger relation with other target-related words while interacting less with the background part.
@@
$\mathcal{S}_{l}=\mathcal{S}_{l-1}+Mamba(LN(\mathcal{S}_{l-1})),l=1,2,...,L$ (9)
@@
In our MiM block, both the inner Mamba block and the outer Mamba block adopt the visual mamba block [28], as shown in Figure  4 (b). The inner Mamba block models the relationship between visual words for local feature extraction, while the outer Mamba block captures the global feature by modelling the relationship between visual sentences. We set the block number $L_{1},L_{2},L_{3},L_{4}=2,4,9,2$ of each stage by default. Stacking the MiM blocks for $L=17$ layers, we build our MiM hierarchical encoder.
@@
3.2.4 Complexity Analysis.
Given a visual sequence $T\in\mathbb{R}^{1\times n\times d}$, the computation complexity of SSM is $8ndN$ [34], where $N$ is the SSM dimension and is set to 16 by default. For simplicity, we only consider the core SSM in SS2D and 3 linear layers and calculate the FlOPs of a visual Mamba block as $8ndN+3nd^{2}$. The MiM block includes an inner Mamba block, an outer Mamba block, and a linear layer, with complexities of $8mncN+3mnc^{2}$, $8ndN+3nd^{2}$, and $nmcd$, where $m$ is the number of visual words in a visual sentence and $c$ is the word embedding dimension. Thus, the total FLOPs for the MiM block $FLOPs_{MiM}$ sum to $8mncN+8ndN+3mnc^{2}+3nd^{2}+nmcd$, still maintaining a linear complexity. While the FLOPs of the standard transformer block $FLOPs_{ST}$ is $2nd(6d+n)$ [30]. Considering that $c\ll d$ and $m\ll n$ in high resolution infrared images, the ratio of $FLOPs_{MiM}$ and $FLOPs_{ST}$ approaches 0, meaning that our MiM block introduces a trivial FLOP increase while offering a superior accuracy-efficiency balance demonstrated in subsequent experiments.
4 Experiments
4.1 Experimental Settings
4.1.1 Datasets.

### v4

# MiM-ISTD: Mamba-in-Mamba for Efficient Infrared Small Target Detection

@@ Abstract
Index Terms:
Mamba-in-Mamba, State Space Model, Infrared Small Target Detection
I Introduction
@@
In recent years, deep-learning-based methods have significantly improved ISTD performances, and most of them are CNN-based methods [12, 13, 14, 15, 16, 17, 10, 18]. However, the drawback of CNN-based methods is that their focus on local features is at the cost of global contexts. Global contexts are also important to ISTD because in infrared images the background pixels and the small targets seem so similar in many cases that cannot be distinguished by local features alone, which easily triggers missed detection. As a solution, some hybrid methods [11, 19, 20, 21, 22, 23, 24] that combine ViT with CNN have been proposed to rely on ViT’s ability to model long-range dependencies. However, these methods generally suffer from a heavier computational burden due to the quadratic computational complexity of ViT. Despite certain work [19] adopting linear ViTs, its accuracy is still subordinate to the designs with quadratic complexity. Considering that high-resolution images are not rare in the infrared remote sensing domain (e.g. images produced by high-resolution infrared military sensors), this efficiency defect will be amplified when the resolution gets larger and hinder real-time ISTD. How to relieve the inefficiency while maintaining high ISTD accuracy is our main concern.
@@
We aim to propose a Mamba-based ISTD encoder to solve the locality defect while still maintaining a superior model efficiency. To this end, we get inspired by TNT [30], which effectively models local structural features with a trivial increase of computation and memory cost, and propose a novel Mamba-in-Mamba (MiM-ISTD) architecture for more efficient ISTD. To boost the feature representation ability of Mamba, we divide the input image into several patches as ”visual sentences” and then further divide them into sub-patches as ”visual words”. We use an Outer Mamba block to extract features of visual sentences, and further assist it with Inner Mamba blocks to excavate the local features of smaller visual words. In particular, features and relations between visual words in each visual sentence are calculated independently using a shared network to ensure the added amount of parameters and FLOPs is minimal. Then, the visual word features are consolidated back into their respective sentences. In this way, MiM-ISTD enables us to extract visual information with refined granularity. We present in Fig. 1 that MiM-ISTD exhibits superior efficiency over other methods in terms of GPU memory usage and inference time, particularly as the resolution of infrared images increases. In general, MiM-ISTD can achieve the most notable accuracy-efficiency balance compared with other SOTA methods.
@@
II Related work
II-A Infrared Small Target Detection Networks
@@
Generally speaking, present ISTD networks can be classified into two categories: CNN-based and hybrid networks. CNN-based networks mainly focus on local feature extraction. Dai et al. [16] propose asymmetric contextual modulation (ACM) for cross-layer information exchange to improve ISTD performance. They also design AlcNet [14], including a local attention module and a cross-layer fusion module to preserve the local features of small targets. Wang et al. propose MDvsFA [17], which applies generative adversarial network (GAN) to ISTD, and achieves a trade-off between miss detection and false alarm. BAUENet [13] introduces uncertainty to ISTD and achieves boundary-aware segmentation. DNANet [10] progressively interacts with high and low-level features. Dim2Clear [31] treats ISTD as an image detail reconstruction task by exploring the image enhancement idea. FC3-Net [32] explores feature compensation and cross-level correlation for ISTD task. ISNet [18] designs a simple Taylor finite difference-inspired block and a two-orientation attention aggregation module to detect targets. Recently, the first diffusion model for ISTD, DCFR-Net[33], has been proposed. However, its accuracy is subordinate to the present SOTA method [11], and its computational efficiency lags considerably when compared to the majority of deep learning-based ISTD techniques.
@@
Relying solely on local features for ISTD may lead to missed detection of small targets, which can merge into similar backgrounds and become indistinguishable. Therefore, hybrid methods complement local details with global contexts by combining ViT with CNN. IRSTFormer[34] adopts hierarchical ViT to model long-range dependencies but lays insufficient emphasis on mining local details. ABMNet [19] adopts ODE methods in both CNN and linear ViT structure design for ISTD. IAANet [20] concatenates local patch outputs from a simple CNN with the original transformer, but causes limited feature extraction, especially in low-contrast scenarios. RKformer [24] applies the Runge-Kutta method to build coupled CNN-Transformer blocks to highlight infrared small targets and suppress background interference. TCI-Former [11] extracts small target features by simulating the thermal conduction process and achieves SOTA results. However, most of these hybrid methods suffer from a quadratic computational complexity due to the usage of ViT. Despite some work [19] adopting linear ViT design, its detection accuracy cannot be on par with other works with quadratic-complexity ViTs.
@@
To improve network efficiency while maintaining high accuracy, we draw inspiration from both Mamba [25] and TNT [30], and propose a Mamba-in-Mamba architecture for ISTD, within which contains a Mamba-in-Mamba (MiM) hierarchical encoder that implements efficient local and global feature extraction.
II-B Mamba in Vision Tasks
@@
Recently, state space sequence models (SSMs) [26] have shown promise in efficiently handling long sequence modeling, offering an alternative for addressing long-range dependencies in visual tasks. Compared with transformers, SSMs are more efficient since they scale linearly with sequence length, and retain a superior ability to model long-range dependencies. Several latest studies have demonstrated the effectiveness of Mamba in vision tasks [35]. For instance, Vim [36] proposed a generic vision backbones with bidirectional Mamba blocks. VMamba [28] proposed a hierarchical Mamba-based vision backbone and a cross-scan module to address the issue of direction-sensitivity arising from the disparities between 1D sequences and 2D image representations.
@@
Notably, Mamba has been most widely applied to the medical image segmentation areas. U-Mamba [27], Vm-unet [37], Mamba-unet [38] and SegMamba [39] proposed a task-specific architecture with the Mamba block based on nnUNet [40], Swin-UNet [41], VMamba [28] and Swin-UNETR [42], respectively. P-Mamba [29] combined PM diffusion with Mamba to efficiently remove background noise while preserving target edge details. Swin-UMamba [43] verified that ImageNet-based pre-training is important to medical image segmentation for Mamba-based networks. Vivim [44] introduced Mamba for medical video object segmentation.
@@
Figure 2: Illustration of the 2D Selective Scan (SS2D) on an infrared image. We commence by scanning an image using scan expanding. The four resulting feature sequences are then individually processed through the S6 block and the four output sequences are merged (scan merging) to construct the final 2D feature map.
@@
Since these Mamba models have achieved promising results in various vision tasks, we intend to study whether Mamba can also bring advancements to ISTD as it has brought to other vision tasks. However, when we directly apply visual Mamba blocks to ISTD, the detection accuracy is not considerable, despite superior model efficiency. The reason is that the targets in ISTD tasks are very small, which requires paying more attention to the local features than other vision tasks with mainly common-size targets, while the original visual Mamba block cannot well explore these local features. To solve this defect, we propose Mamba-in-Mamba (MiM-ISTD) for ISTD, which takes visual sentences and visual words flows simultaneously and sends each flow to respective visual Mamba blocks to obtain both local and global features with high efficiency.
III Proposed method
@@
III-A Preliminaries
III-A1 State Space Models.
State Space Models (SSMs) are commonly employed as linear time-invariant systems that transform a one-dimensional input stimulus $x(t)\in\mathbb{R}^{L}$ through intermediary implicit states $h(t)\in\mathbb{R}^{N}$ to an output $y(t)\in\mathbb{R}^{L}$. In mathematical terms, SSMs are typically described by linear ordinary differential equations (ODEs) ((1)), where the system is characterized by a set of parameters including the state transition matrix $A\in\mathbb{C}^{N\times N}$, the projection parameters $B,C\in\mathbb{C}^{N}$, and the skip connection $D\in\mathbb{C}^{1}$.
@@ III-A Preliminaries
III-A2 Discretization.
State Space Models (SSMs) present significant challenges when applied to deep learning scenarios due to their continuous-time nature. To address this, the ODE needs to be transformed into a discrete function. Considering the input $x_{k}\in\mathbb{R}^{L\times D}$, a sampled vector within the signal flow of length $L$ following [45], a timescale parameter $\Delta$ can be introduced to the continuous parameters $A$ and $B$ into their discrete counterparts $\overline{A}$ and $\overline{B}$ following the zeroth-order hold (ZOH) rule. Consequently, (1) can be discretized as follows:
$\begin{split}&h_{k}=\overline{A}h_{k-1}+\overline{B}x_{k},\quad y(t)=Ch(t)+Dx(t)\\ &\overline{A}=e^{\Delta A},\quad\overline{B}=(e^{\Delta A}-I)A^{-1}B,\quad\overline{C}=C\\ \end{split}$ (2)
@@
III-A3 2D Selective Scan.
2D spatial information cannot be effectively captured by models designed for 1D data, making it unsuitable to directly apply Mamba to vision tasks. The 2D selective scan (SS2D) in [28] can address the issue. The overview of SS2D is depicted in (4). SS2D arranges image patches in four different directions to generate four separate sequences. The quad-directional scanning strategy ensures that each element in the feature map integrates information from all other locations in different directions, thus creating a global receptive field without increasing linear computational complexity. Each resulting feature sequence is then processed using the selective scan space state sequential model (S6) before merging to reconstruct the 2D feature map. Given the input feature $z$, then the output feature $\overline{z}$ of SS2D can be expressed as:
@@
where $i\in\{1,2,3,4\}$ represents one of the four scanning directions. $expand(·)$ and $merge(·)$ refer to the scan expanding and scan merging operations in [28]. The S6 block in (4) enables each element in a 1D array to engage with any previously scanned samples through a condensed hidden state. For a more comprehensive understanding of S6, [28] provides an in-depth explanation.
@@
III-B Mamba-in-Mamba for Efficient ISTD
Present ISTD methods mainly typically employ CNNs or hybrids of CNNs and ViTs. The latter compensates for the shortcomings of the former in modeling long-range dependencies but suffers from a quadratic computational complexity. Recently, Mamba has been proposed. It is renowned for its superior model efficiency, less GPU memory usage, and better long-range dependency modeling and has been successfully applied to the vision domain. Therefore, we explore whether Mamba can also be applied to improve ISTD performances. However, when we directly apply visual mamba block to ISTD, the accuracy is not very impressive because local features, which matter a lot to detect small targets, are less explored. To address this deficiency, we propose a Mamba-in-Mamba (MiM-ISTD) architecture, shown in Fig. 3, to learn both global and local information in an image while guaranteeing superb model efficiency.
@@ III-B Mamba-in-Mamba for Efficient ISTD
III-B1 Convolutional Stem.
We construct a convolutional stem, where a stack of $3\times 3$ convolutions is utilized, to produce visual words $\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$ and visual sentences $\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times D}$ at the first stage, where $C$ is the visual word dimension and $D$ is the visual sentence dimension. Each visual word corresponds to a $2\times 2$ pixel region in the original image, and each visual sentence is composed of $4\times 4$ visual words. Unlike ViTs, we do not add the position embedding bias to visual words and sentences due to the causal nature of visual mamba block [28].
Figure 4: Overview of (a) our proposed Mamba-in-Mamba (MiM) block, which contains an Inner Mamba block and an Outer Mamba block, and (b) the structure of Inner/Outer Mamba block [28]. The Inner Mamba block is shared in the same layer. The dashed line in (a) means bypassing the Outer Mamba block.
III-B2 MiM Hierarchical Encoder.
The core part of our MiM is its hierarchical encoder of four stages with different numbers of tokens, as shown in Fig. 3. Across the four stages, the spatial shape of visual words is set as $\frac{H}{2}\times\frac{W}{2}$, $\frac{H}{4}\times\frac{W}{4}$, $\frac{H}{8}\times\frac{W}{8}$ and $\frac{H}{16}\times\frac{W}{16}$. The spatial shape of visual sentences are set as $\frac{H}{8}\times\frac{W}{8}$, $\frac{H}{16}\times\frac{W}{16}$, $\frac{H}{32}\times\frac{W}{32}$ and $\frac{H}{64}\times\frac{W}{64}$. We adopt the patch merging in [41] as the down-sampling operation. Each stage consists of multiple MiM blocks which process both word-level and sentence-level features. The visual words $x^{i,j}$ are mapped to a sequence of word embeddings $w^{i,j}$ via a linear projection, expressed as:
@@
The MiM block handles two different data streams: one traverses through the visual sentences, while the other manages the visual words within each sentence. For the word embeddings, the relation between visual words should be exploited as follows:
$\begin{split}&W^{i^{\prime}}_{l}=W^{i}_{l-1}+Mamba(LN(W^{i}_{l-1}))\\ &W^{i}_{l}=W^{i^{\prime}}_{l}+Conv_{FFN}(LN(W^{i^{\prime}}_{l}))\\ \end{split}$ (7)
where $l$ represents the index corresponding to the $l^{th}$ block within a series spanning from $1$ to $L$, with $L$ denoting the aggregate number of such blocks. The first word embedding input $W^{i}_{0}$ is the $W^{i}$ in (6). $W^{i^{\prime}}_{l}$ denotes the intermediate feature. All transformed word embeddings are denoted by $\mathcal{W}_{l}=[W^{1}_{l},W^{2}_{l},...,W^{n}_{l}]$. This can be viewed as an Inner Mamba block, illustrated in Fig. 4 (a). In this process, the relationships among visual words are built by computing interactions among each two visual words. For example, in a patch containing a small target, a word denoting the target would have a stronger relation with other target-related words while interacting less with the background part.
@@
$\begin{split}&\mathcal{S}^{{}^{\prime}}_{l}=\mathcal{S}_{l-1}+Mamba(LN(\mathcal{S}_{l-1}))\\ &\mathcal{S}_{l}=\mathcal{S}^{{}^{\prime}}_{l}+Conv_{FFN}(LN(\mathcal{S}^{{}^{\prime}}_{l}))\\ \end{split}$ (9)
@@
Within our MiM block, the visual mamba block [28] is employed in both the Inner and Outer Mamba blocks, each followed by a coupling with a convolutional feed-forward network, as shown in Fig. 4 (b). The Inner Mamba block models the relationship between visual words for local feature extraction, while the Outer Mamba block captures the global feature by modelling the relationship between visual sentences. The convolutional feed-forward network is responsible for augmenting these features with finer local details. We set the block number $L_{1},L_{2},L_{3},L_{4}=2,2,2,2$ of each stage by default. Stacking the MiM blocks for $L=8$ layers, we build our MiM hierarchical encoder.
@@
III-B4 Complexity Analysis.
Given a visual sequence $T\in\mathbb{R}^{1\times n\times d}$, the computation complexity of SSM is $3nEN+nEN=96nd+32nd=128nd$ [36], where $N$ is the SSM dimension and is set to 16 by default and $E=2d$. In comparison, the computation complexity of self-attention is $4nd^{2}+2n^{2}d$. It is evident that self-attention exhibits quadratic complexity to the sequence length $n$, whereas SSM is linear. This computational efficiency makes our MiM-ISTD more scalable compared to other quadratic Transformer-based models like IAANet [20] and RKformer [24].
Each of our proposed MiM block includes 16 Inner Mamba block and an Outer Mamba block, with each Mamba block containing the core visual Mamba part and a convolutional feed forward network. For simplicity, we only consider the most critical visual Mamba when calculating FLOPs. This integral part encompasses an SSM and three linear layers. We compute its FLOPs as $128nd+3nd^{2}$. Therefore, the FLOPs of the several Inner Mamba blocks in all and an Outer Mamba block in an MiM block can be calculated as $128mnc+3mnc^{2}$ and $128nd+3nd^{2}$, where $m$ is the number of visual words in a visual sentence and $c$ is the word embedding dimension. Thus, the total FLOPs for the MiM block $FLOPs_{MiM}$ sum to $128mnc+128nd+3mnc^{2}+3nd^{2}$, still maintaining a linear complexity. While the FLOPs of the standard transformer block $FLOPs_{ST}$ is $2nd(6d+n)$ [30]. Considering that $c\ll d$ and $m\ll n$ in high resolution infrared images, the ratio of $FLOPs_{MiM}$ and $FLOPs_{ST}$ approaches 0, meaning that our MiM block introduces a trivial FLOP increase while offering a superior accuracy-efficiency balance demonstrated in subsequent experiments.
IV Experiments
IV-A Experimental Settings
IV-A1 Datasets.